# Nearest Neighbors GParareal: Improving Scalability of Gaussian Processes for Parallel-in-Time Solvers

### v1

@@ 1 Introduction
In this paper, we focus on improving the scalability performance of a particular PinT solver, GParareal [44], reducing its computational cost while maintaining the same accuracy. GParareal is built upon Parareal [33], an iterative algorithm that uses two solvers of different granularity: a fine one, denoted by $\mathscr{F}$, assumed accurate but slow, and a coarse one, denoted by $\mathscr{G}$, faster but less precise. First, the time span is divided into $N$ subintervals and the $\mathscr{G}$ solver is run sequentially on one core. Then, the computationally expensive numerical integrator $\mathscr{F}$ is run in parallel on $N$ cores, one for each time interval, starting from the initial condition provided by $\mathscr{G}$. All initial conditions are then iteratively updated using an evaluation of $\mathscr{G}$ and a correction term based on the discrepancy $\mathscr{F}-\mathscr{G}$ between the two solvers computed at the previous iteration. Parareal stops once a desired tolerance is met, with the convergence of its solution to the true one guaranteed under certain mathematical conditions (see Section 2). The algorithm proceeds in a Markovian manner, relying only on the last iteration’s data to propagate the algorithm further. This is potentially wasteful, as previously generated data can be leveraged within a Machine Learning framework. In particular, neural networks and their physics-informed versions (PINNs) [8] have been used to learn and approximate different components of the procedure, such as the fine solver $\mathscr{F}$, reducing the running time, or the course solver $\mathscr{G}$, aiming to improve the course approximation, favoring a faster convergence (see, e.g., [1, 28, 37, 41, 61]). However, many of these studies precisely state neither the number of used processors, nor the computational cost of neural networks’ training, as the latter is often assumed to be “offline” and thus not included in Parareal’s runtimes.
@@ 3 GParareal
While Parareal leads to speed-up and more efficient use of resources, generally achieving convergence in $K_{\rm Para}<N$ iterations, $K_{\rm Para}$ can still be a sensible fraction of $N$, which leaves room for improvements (see Table 3). GParareal was proposed for this [44]. Instead of computing the predictor-corrector (2) at iteration $k$ using information from the previous iteration $k-1$, as done in Parareal, GParareal uses data from the current iteration $k$, using the following expression
@@
with posterior mean $\mu_{\mathcal{D}_{k}}^{(s)}(\boldsymbol{U}^{\prime})\in\mathbb{R}$ and variance $\mathcal{K}_{\mathcal{D}_{k}}(\boldsymbol{U}^{\prime},\boldsymbol{U}^{\prime})\in\mathbb{R}$ given by
@@
where $\mathcal{K}({U},\boldsymbol{U}^{\prime})\in\mathbb{R}^{Nk}$ is a vector of covariances between every input collected in ${U}$ and $\boldsymbol{U}^{\prime}$ defined as $(\mathcal{K}({U},\boldsymbol{U}^{\prime}))_{r}=\mathcal{K}_{\rm GP}({{U}_{(r,\cdot)}}^{\top},\boldsymbol{U}^{\prime})$, $r=1,\ldots,Nk$, and $\mathcal{K}({U},{U})\in\mathbb{R}^{Nk\times Nk}$ is the covariance matrix, with $(\mathcal{K}({U},{U}))_{r,q}=\mathcal{K}_{\rm GP}({{U}_{(r,\cdot)}}^{\top},{{U}_{(q,\cdot)}}^{\top})$, $r,q=1,\ldots,Nk$. Note that (10) slightly differs from the posterior variance in [44], as we added a regularization term $\sigma_{\rm reg}^{2}$, known as nugget, jitter, or regularization strength, to improve the condition number of the covariance matrix upon inversion (see Supplementary Material G for the detailed discussion).
An important step in the training of the GP is the optimization of the hyperparameters $\boldsymbol{\theta}:=(\sigma_{\rm i}^{2},\sigma_{\rm o}^{2},\sigma_{\rm reg}^{2})$ for each of the $d$ components, which is done by numerically maximizing the marginal log-likelihood (see Supplementary Material G for implementation details):
@@
where $\mathcal{K}(\cdot,\cdot)$ depends on $\boldsymbol{\theta}$ through the kernel $\mathcal{K}_{\rm GP}$ in (6). Once the optimal hyper-parameters $\boldsymbol{\theta}^{*}$ are determined, the $d$ scalar posterior distributions of the GP (8) can be used to make a prediction at $\boldsymbol{U}^{\prime}$.
@@ 4.1 Rethinking the predictor-corrector rule
A special case is when $m=1$, for which the prediction equals the output of the closest point.
@@ 4.5 Computational complexity
Let $T_{\mathscr{G}}$ and $T_{\mathscr{F}}$ be the time it takes to run $\mathscr{G}$ and $\mathscr{F}$ over one interval $[t_{i},t_{i+1}]$, respectively. Here, we compute the parallel speed-up $S_{\rm alg}\in\mathbb{R}^{+}$ of using the Parareal, GParareal, and nnGParareal algorithms as opposed to running $\mathscr{F}$ serially across $N$ intervals. The speed-up describes the parallel efficiency of the algorithm and is computed as the ratio of the serial runtime $T_{\textrm{Serial}}:=NT_{\mathscr{F}}$ over the parallel runtime $T_{\rm alg}$, that is $S_{\rm alg}:=T_{\textrm{Serial}}/T_{\rm alg}$. $S_{\rm alg}$ can be approximated theoretically, as done below, or computed empirically, as done in Section 5. We refer to the latter as $\hat{S}_{\rm alg}$. Let $T_{\textrm{model}}(k)$ be the total cost of evaluating $\widehat{f}_{\textrm{model}}$ during iteration $k$, which includes both model training and predicting. If an algorithm converges in $K_{\rm alg}$ iterations, its worst-case runtime, excluding serial and parallel overheads, can be approximated by
@@
For Parareal, $T_{\textrm{model}}=T_{\rm Para}\in O(1)$, since $\widehat{f}_{\rm Para}$ is a simple look-up operation. This would be the fastest parallel algorithm to converge if the three algorithms were converging at the same iteration $K$, which is typically not the case. For the GP, $T_{\textrm{model}}(k)=T_{\rm GP}(k)$ is the wallclock time needed to run the $d$ independent GPs in parallel at iteration $k$, which includes fitting the model, optimizing the hyperparameters, and making predictions. This quantity is challenging to estimate beforehand, as it depends on the size of the covariance matrix to invert and the iterations required to optimize the log-likelihood when computing the optimal hyperparameters. Nonetheless, since the size of the dataset $\mathcal{D}_{k}$ for one coordinate is of order $O(kN)$ by iteration $k$, $T_{\rm GP}(k)$ scales as $O((d/N\vee 1)k^{3}N^{3})$ [44], where $(d/N\vee 1)$ is implied by the fact that the $d$ GPs are independent and can be parallelized over the $N$ cores. Overall, we have
$T_{\rm GP}=\sum_{k=1}^{K_{\rm GPara}}T_{\rm GP}(k)\in O\left((d/N\vee 1)K_{\rm GPara}^{4}N^{3}\right).$
@@
Unlike GParareal, the runtime $T_{{\rm nnGPara}}$ of nnGParareal can be accurately estimated as the reduced subset is of constant size $m$. Let $n_{\rm start}$ be the number of random restarts performed during the optimization procedure, $n_{\rm reg}$ the numbers of tested jitters values $\sigma^{2}_{\rm reg}$ (see also Supplementary Material G), and $T_{{\rm GP}_{m}}$ the cost of training, hyperparameter optimization, and prediction for a GP based on a dataset of size $m$. Then, the model cost of nnGParareal at iteration $k$ is
$T_{\rm nnGP}(k)=(N-k)\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)T_{{\rm GP}_{m}}=(N-k)T_{\rm nnGP}^{m},$
where the rescaling term $(N-k)$ follows from the need to retrain the nnGP at every unconverged interval at iteration $k$, while $T_{\rm nnGP}^{m}$ is the cost of using the model to make a single prediction, including training. Unlike GParareal, $T_{\rm nnGP}(k)$ decreases as $k$ increases, see Figure 5. Similarly to GParareal, the quantity $dn_{\rm reg}n_{\rm start}$ can be massively parallelized over the $N$ cores. Moreover, $T_{\rm nnGP}^{m}$ is almost identical across intervals and iterations. Hence, to estimate $T_{\rm nnGPara}$, it is sufficient to estimate $T_{{\rm GP}_{m}}$, which is straightforward given the small sample size required. Although the advantage of using nnGParareal may not immediately clear from the above formula, $T_{\rm nnGP}^{m}$ is sensibly cheaper than $T_{\rm GP}$, as
$T_{\rm nnGP}^{m}\in O\left(\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)(m^{3}+\log(kN))\right).$
The term $m^{3}$ comes from the cubic cost of inverting a covariance matrix of size $m$, while $\log(kN)$ provides the cost of determining the $m$ nearest neighbors. Among several approaches developed for the latter, we opt for kd-trees [18] for structured storage of the observations and fast retrieval of the closest points. Searching the $m$ nns has an average complexity of $O(\log N)$. Updating the tree structure for the whole iteration $k$ has a total cost of $O(N\log(kN))$, since the insertion of a tree element is in $O(\log n)$, with the number $n$ of observations currently in the tree, and the cost for finding $1$ or $m$ neighbors is the same (see [18] or [55] for a summary of kd-tree operations cost). Hence, $T_{\rm nnGP}=\sum_{k=1}^{K_{{\rm nnGPara}}}T_{\rm nnGP}(k)$ becomes
$T_{\rm nnGP}\in O(\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)K_{{\rm nnGPara}}N(m^{3}+\log(K_{{\rm nnGPara}}N)),$
which, for small values of $m$ (e.g., $15-20$), is much lower than $O(K_{\rm GPara}^{4}N^{3})$ for $T_{\rm GP}$.
@@ 4.6 Error Analysis
In this section, we provide an upper bound for the error between the exact and the nnGParareal solutions at iteration $k$ and interval $i$. A similar bound has been provided for GParareal [44, Theorem 3.5], which relies on the posterior mean consistency of GPs. There, the GP error is controlled by the fill distance, defined as
@@
$=\|\mathscr{F}(\boldsymbol{U}_{i-1})-\mathscr{G}(\boldsymbol{U}_{i-1}^{k+1})-\widehat{f}(\boldsymbol{U}_{i-1}^{k+1})\pm f(\boldsymbol{U}_{i-1}^{k+1})\pm\mathscr{G}(\boldsymbol{U}_{i-1})\|.$
By using the triangle inequality and the definition of $f$, we obtain
@@
(20) $E_{i}^{k+1}\leq\beta E_{i-1}^{k+1}+\alpha E_{i-1}^{k+1}+\eta(E_{i-1}^{k}+E_{i-1}^{k+1})=(\alpha+\beta+\eta)E_{i-1}^{k+1}+\eta E_{i-1}^{k}.$
@@
$E_{i}^{0}$ $=\|\boldsymbol{U}_{i}-\boldsymbol{U}_{i}^{0}\pm g(\boldsymbol{U}_{i-1})\|$
$\leq\|\mathscr{G}(\boldsymbol{U}_{i-1})-\mathscr{G}(\boldsymbol{U}_{i-1}^{0})\|+\|\boldsymbol{U}_{i}-\mathscr{G}(\boldsymbol{U}_{i-1})\|=\beta E^{0}_{i-1}+\gamma\leq(\alpha+\beta+\eta)E_{i-1}^{0}+\gamma,$
$i\in\mathbb{N}$, where Assumptions 1-3 are used and $\gamma=C_{3}\Delta t^{p+1}$. This recursive inequality can be solved using the result from Lemma B.1 in [45], with their constants $\Lambda=0,A=\eta,B=(\alpha+\beta+\eta),D=\gamma,n=i$ and adjusting for the initial error being $E_{i}^{0}$ instead of $E_{i}^{1}$.
@@ 5.2 Non-linear Hopf bifurcation model
(21) $\frac{du_{1}}{dt}=-u_{2}+u_{1}\left(\frac{t}{T}-u_{1}^{2}-u_{2}^{2}\right),\quad\frac{du_{2}}{dt}=u_{1}+u_{2}\left(\frac{t}{T}-u_{1}^{2}-u_{2}^{2}\right).$

### v2

@@ 1 Introduction
In this paper, we focus on improving the scalability performance of a particular PinT solver, GParareal [45], reducing its computational cost while maintaining the same accuracy. GParareal is built upon Parareal [34], an iterative algorithm that uses two solvers of different granularity: a fine one, denoted by $\mathscr{F}$, assumed accurate but slow, and a coarse one, denoted by $\mathscr{G}$, faster but less precise. First, the time span is divided into $N$ subintervals and the $\mathscr{G}$ solver is run sequentially on one core. Then, the computationally expensive numerical integrator $\mathscr{F}$ is run in parallel on $N$ cores, one for each time interval, starting from the initial condition provided by $\mathscr{G}$. All initial conditions are then iteratively updated using an evaluation of $\mathscr{G}$ and a correction term based on the discrepancy $\mathscr{F}-\mathscr{G}$ between the two solvers computed at the previous iteration. Parareal stops once a desired tolerance is met, with the convergence of its solution to the true one obtained by running $\mathcal{F}$ serially guaranteed under certain mathematical conditions (see Section 2). The algorithm proceeds in a Markovian manner, relying only on the last iteration’s data to propagate the algorithm further. This is potentially wasteful, as previously generated data can be leveraged within a Machine Learning framework. In particular, neural networks and their physics-informed versions (PINNs) [8] have been used to learn and approximate different components of the procedure, such as the fine solver $\mathscr{F}$, reducing the running time, or the course solver $\mathscr{G}$, aiming to improve the course approximation, favoring a faster convergence (see, e.g., [1, 29, 42, 62]). Recently, Parareal has also been used to speed-up the training of PINNs in [38].
However, many of these studies precisely state neither the number of used processors, nor the computational cost of neural networks’ training, as the latter is often assumed to be “offline” and thus not included in Parareal’s runtimes.
@@ 3 GParareal
While Parareal leads to speed-up and more efficient use of resources with respect to running the fine solver sequentially, achieving convergence in $K_{\rm Para}\leq N$ iterations, $K_{\rm Para}$ can still be a sensible fraction of $N$, which leaves room for improvements (see Table 3). GParareal was proposed for this [45]. Instead of computing the predictor-corrector (2) at iteration $k$ using information from the previous iteration $k-1$, as done in Parareal, GParareal uses data from the current iteration $k$, using the following expression
@@
with posterior mean $\mu_{\mathcal{D}_{k}}^{(s)}(\boldsymbol{U}^{\prime})\in\mathbb{R}$ and variance $\mathcal{K}_{\mathcal{D}_{k}}(\boldsymbol{U}^{\prime},\boldsymbol{U}^{\prime})\in\mathbb{R}^{+}$ given by
@@
where $\mathcal{K}({U},\boldsymbol{U}^{\prime})\in\mathbb{R}^{Nk}$ is a vector of covariances between every input collected in ${U}$ and $\boldsymbol{U}^{\prime}$ defined as $(\mathcal{K}({U},\boldsymbol{U}^{\prime}))_{r}=\mathcal{K}_{\rm GP}({{U}_{(r,\cdot)}}^{\top},\boldsymbol{U}^{\prime})$, $r=1,\ldots,Nk$, and $\mathcal{K}({U},{U})\in\mathbb{R}^{Nk\times Nk}$ is the covariance matrix, with $(\mathcal{K}({U},{U}))_{r,q}=\mathcal{K}_{\rm GP}({{U}_{(r,\cdot)}}^{\top},{{U}_{(q,\cdot)}}^{\top})$, $r,q=1,\ldots,Nk$. Note that (10) slightly differs from the posterior variance in [45], as we added a regularization term $\sigma_{\rm reg}^{2}$, known as nugget, jitter, or regularization strength, to improve the condition number of the covariance matrix upon inversion (see Supplementary Material I for the detailed discussion).
An important step in the training of the GP is the optimization of the hyperparameters $\boldsymbol{\theta}:=(\sigma_{\rm i}^{2},\sigma_{\rm o}^{2},\sigma_{\rm reg}^{2})$ for each of the $d$ components, which is done at every iteration by numerically maximizing the marginal log-likelihood (see Supplementary Material I for implementation details):
@@
where $\mathcal{K}(\cdot,\cdot)$ depends on $\boldsymbol{\theta}$ through the kernel $\mathcal{K}_{\rm GP}$ in (6). Once the optimal hyperparameters $\boldsymbol{\theta}^{*}$ are determined, the $d$ scalar posterior distributions of the GP (8) can be used to make a prediction at $\boldsymbol{U}^{\prime}$.
@@ 4.1 Rethinking the predictor-corrector rule
A special case is when $m=1$, for which the prediction equals the output of the closest point. While the Euclidian distance is a natural choice, alternatives like Minkowski or Mahalanobis distances may also be used. However, our findings (not shown) indicate that the choice of distance has a marginal impact on performance. Therefore, we consider the Euclidean distance in this work.
@@ 4.5 Computational complexity
Let $T_{\mathscr{G}}$ and $T_{\mathscr{F}}$ be the time it takes to run $\mathscr{G}$ and $\mathscr{F}$ over one interval $[t_{i},t_{i+1}]$, respectively. Here, we compute the parallel speed-up $S_{\rm alg}\in\mathbb{R}^{+}$ of using the Parareal, GParareal, and nnGParareal algorithms as opposed to running $\mathscr{F}$ serially across $N$ intervals. The speed-up describes the relative performance of two systems processing the same problem and is computed as the ratio of the serial runtime $T_{\textrm{Serial}}:=NT_{\mathscr{F}}$ over the parallel runtime $T_{\rm alg}$, that is $S_{\rm alg}:=T_{\textrm{Serial}}/T_{\rm alg}$. $S_{\rm alg}$ can be approximated theoretically, as done below, or computed empirically (from measured runtimes including overheads), as done in Section 5. We refer to the latter as $\hat{S}_{\rm alg}$. Let $T_{\textrm{model}}(k)$ be the total cost of evaluating $\widehat{f}_{\textrm{model}}$ during iteration $k$, which includes both model training and predicting. If an algorithm converges in $K_{\rm alg}$ iterations, its worst-case runtime, excluding communication overheads, can be approximated by
@@
For Parareal, $T_{\textrm{model}}=T_{\rm Para}\in O(1)$, since $\widehat{f}_{\rm Para}$ is a simple look-up operation. This would be the fastest parallel algorithm to converge if the three algorithms were converging at the same iteration $K$, which is typically not the case. For the GP, $T_{\textrm{model}}(k)=T_{\rm GP}(k)$ is the wallclock time needed to run the $d$ independent GPs in parallel at iteration $k$, which includes fitting the model, optimizing the hyperparameters, and making predictions. This quantity is challenging to estimate beforehand, as it depends on the size of the covariance matrix to invert and the iterations required to optimize the log-likelihood when computing the optimal hyperparameters. Nonetheless, since the size of the dataset $\mathcal{D}_{k}$ for one coordinate is of order $O(kN)$ by iteration $k$, and $(N-k)$ predictions are made, $T_{\rm GP}(k)$ scales as
$\underbrace{(\tfrac{d}{N}\vee 1)\;O\left(d(Nk)^{2}+(Nk)^{3}\right)}_{\text{Training}}+\underbrace{(\tfrac{d}{N}\vee 1)\;(N-k)\;O(dNk)}_{\text{Post. mean prediction}},$
where $(d/N\vee 1)$ is implied by the fact that the $d$ GPs are independent and can be parallelized over the $N$ cores, $d(Nk)^{2}$ corresponds to computing the kernel matrix $\mathcal{K}({X},{X})$, $(Nk)^{3}$ for its inversion, and $O(dNk)$ is the cost of one posterior mean evaluation. Overall, we have
$T_{\rm GP}=\sum_{k=1}^{K_{\rm GPara}}T_{\rm GP}(k)\in O\left((d/N\vee 1)(d\vee Nk)K_{\rm GPara}^{3}N^{2}\right).$
@@
Unlike GParareal, the runtime $T_{{\rm nnGPara}}$ of nnGParareal can be accurately estimated as the reduced subset is of constant size $m$. Let $n_{\rm start}$ be the number of random restarts performed during the optimization procedure, $n_{\rm reg}$ the numbers of tested jitters values $\sigma^{2}_{\rm reg}$ (see also Supplementary Material I), and $T_{{\rm GP}_{m}}$ the cost of training, hyperparameter optimization, and prediction for a GP based on a dataset of size $m$. Then, the model cost of nnGParareal at iteration $k$ is
$T_{\rm nnGP}(k)=(N-k)\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)T_{{\rm GP}_{m}}=(N-k)T_{\rm m-nnGP},$
where the rescaling term $(N-k)$ follows from the need to retrain the nnGP at every unconverged interval at iteration $k$, while $T_{\rm m-nnGP}$ is the cost of using the model to make a single prediction, including training. Unlike GParareal, $T_{\rm nnGP}(k)$ decreases as $k$ increases, see Figure 5. Similarly to GParareal, the quantity $dn_{\rm reg}n_{\rm start}$ can be massively parallelized over the $N$ cores. Moreover, $T_{\rm m-nnGP}$ is almost identical across intervals and iterations. Hence, to estimate $T_{\rm nnGPara}$, it is sufficient to estimate $T_{{\rm GP}_{m}}$, which is straightforward given the small sample size required. Although the advantage of using nnGParareal may not immediately clear from the above formula, $T_{\rm m-nnGP}$ is sensibly cheaper than $T_{\rm GP}$, as
$T_{\rm m-nnGP}\in O\left(\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)((d\vee m)m^{2}+\log(kN))\right).$
The term $(d\vee m)m^{2}$ comes from the cost of evaluating and inverting a covariance matrix of size $m$, while $\log(kN)$ provides the cost of determining the $m$ nearest neighbors. Among several approaches developed for the latter, we opt for kd-trees [18] for structured storage of the observations and fast retrieval of the closest points. Searching the $m$ nns has an average complexity of $O(\log N)$. Updating the tree structure for the whole iteration $k$ has a total cost of $O(N\log(kN))$, since the insertion of a tree element is in $O(\log n)$, with the number $n$ of observations currently in the tree, and the cost for finding $1$ or $m$ neighbors is the same (see [18] or [56] for a summary of kd-tree operations cost). Hence, $T_{\rm nnGP}=\sum_{k=1}^{K_{{\rm nnGPara}}}T_{\rm nnGP}(k)$ becomes
$T_{\rm nnGP}\in O(\left((dn_{\rm reg}n_{\rm start})/N\vee 1\right)K_{{\rm nnGPara}}N((d\vee m)m^{2}+\log(K_{{\rm nnGPara}}N)),$
which, for small values of $m$ (e.g., $15-20$), is much lower than that of $T_{\rm GP}$.
@@ 4.6 Error Analysis
In this section, we provide an upper bound for the error between the exact and the nnGParareal solutions at iteration $k$ and interval $i$. A similar bound has been provided for GParareal [45, Theorem 3.5], which relies on the posterior mean consistency of GPs. There, the GP error is controlled by the fill distance, defined as
@@
$=\|\mathscr{F}(\boldsymbol{U}_{i-1})-\mathscr{G}(\boldsymbol{U}_{i-1}^{k+1})-\widehat{f}(\boldsymbol{U}_{i-1}^{k+1})\pm f(\boldsymbol{U}_{i-1}^{k+1})\pm\mathscr{G}(\boldsymbol{U}_{i-1})\|,$
where $\pm$ denotes adding and subtracting the corresponding term. By using the triangle inequality and the definition of $f$, we obtain
@@
$E_{i}^{k+1}\leq\beta E_{i-1}^{k+1}+\alpha E_{i-1}^{k+1}+\eta(E_{i-1}^{k}+E_{i-1}^{k+1})=(\alpha+\beta+\eta)E_{i-1}^{k+1}+\eta E_{i-1}^{k}.$
@@
$E_{i}^{0}$ $=\|\boldsymbol{U}_{i}-\boldsymbol{U}_{i}^{0}+\mathscr{G}(\boldsymbol{U}_{i-1})-\mathscr{G}(\boldsymbol{U}_{i-1})\|\leq\|\mathscr{G}(\boldsymbol{U}_{i-1})-\mathscr{G}(\boldsymbol{U}_{i-1}^{0})\|+\|\boldsymbol{U}_{i}-\mathscr{G}(\boldsymbol{U}_{i-1})\|$
$=\beta E^{0}_{i-1}+\gamma\leq(\alpha+\beta+\eta)E_{i-1}^{0}+\gamma,\quad i\in\mathbb{N},$
where Assumptions 1-3 are used and $\gamma=C_{3}\Delta t^{p+1}$. This recursive inequality can be solved using the result from Lemma B.1 in [46], with their constants $\Lambda=0,A=\eta,B=(\alpha+\beta+\eta),D=\gamma,n=i$ and adjusting for the initial error being $E_{i}^{0}$ instead of $E_{i}^{1}$.
@@ 5.2 Non-linear Hopf bifurcation model
$\frac{du_{1}}{dt}=-u_{2}+u_{1}\left(\frac{t}{T}-u_{1}^{2}-u_{2}^{2}\right),\quad\frac{du_{2}}{dt}=u_{1}+u_{2}\left(\frac{t}{T}-u_{1}^{2}-u_{2}^{2}\right).$
@@
Finally, we investigate the impact of the coarse solver accuracy (measured as the number of coarse solvers steps) on the Parareal and nnGParareal speed-ups, for $N=32$ and $1.74e^{8}$ steps of the fine solver. The results (see Figure 15 in Supplementary Material D) show that nnGParareal can use a coarser solver than Parareal to obtain a similar speed-up, with the latter showing non-convergent behavior for smaller numbers of coarse solver steps.